November 19, 2013: It appears that the pump that feeds water into the seakeeper box is not functioning properly. All water property data are affected including temperature, conductivity, salinity, chlorophyll, turbidity, fluorescence and oxygen. All data have been flagged accordingly and an update will be posted once the problem is resolved. A maintenance visit is scheduled for early December that will hopefully resolve the issue – an update will be posted.
UPDATE November 19, 2013: The Q of A is back in service. Data from the Seakeeper system are available beginning November 16, 2013. The replaced devices include the TSG, Wetlabs Eco Triplet and Aanderra Optode. See the latest maintenance report for details. Note that it appears that we are experiencing some flow issues – these data will be affected – we are monitoring to determine the magnitude of the problem.
The Q of A will be temporarily out of service from October 16-November 15th, 2013. No data will be reported during this time period. For more information, please see the notice on the BC Ferries website here.
There was an issue with connecting to the DDL which caused the ports to be turned off. VENUS engineers have diagnosed and fixed the error – we are currently monitoring the data to ensure that it does not happen again.
UPDATE November 19, 2013: The SOG Meteorological Station was replaced and data are available as of November 9. See here for plots.
As of October 28th at approximately 2130 UTC, the SOG Meteorological Station went down and we are not able to communicate with it. The station will be undergoing servicing soon and data acquisition and access will be restored. An update will be posted when the station is up and running.
The ONC Digital Infrastructure team will be conducting a software release between 13:00-14:00 PST today (Thursday November 14 2013). All ONC data services will be interrupted during this time. There will be small gaps in data streams from Saanich Inlet and Strait of Georgia.
The Wetlabs NTUS 319 located on the Microsquid/Digital Camera Frame has lost communication. The last reported data from this device are at ~21:41:39 on November 9, 2013.
UVic Systems’ MATLAB license server went down yesterday evening so many data requests from the ONC data access interfaces (Oceans 2.0 Data Search and VENUS Download Data) failed. The server has been restored and ONC data services are functioning again.
Our LDAP server also failed yesterday evening. So users could not login to access any ONC data between 2013-Nov-06 ~5:00pm and ~9:00am (PST). Our Systems team resolved the issue early this morning.
All data requests from VENUS Download Data are failing and our Data Plots aren’t updating. The ONC Systems team is currently investigating the problem. We hope to have the issue resolved as soon as possible.
UPDATE: 2013-Nov-06 11:20 PST
The problem has been resolved. An overloaded web server was causing the problem. The Systems team rebooted it and now data can now be requested from the VENUS Download Data page and Data Plots are updating.
There is an issue with all our ASL Zooplankton Acoustic Profiler (ZAP) data products. We are working to resolve the problem, but right now all the data products available from the VENUS Download Data are producing errors.
UDPATE: 2013-Nov=06 11:25 PST
The ONC Systems team rebooted one of the web servers which has resolved the problem. ZAP data products can now be requested from the VENUS Download Data page.
The VENUS web site is undergoing maintenance from Oct 22-25, 2013. No outages are expected.